OpenStudy (anonymous):

differentiate -5e^4x

zepdrix (zepdrix):

The constant won't affect the differentiation process, so we can pull it out front. \[\Large \color{royalblue}{\left(\color{black}{-5e^{4x}}\right)'}\quad=\quad\color{black}{-5}\color{royalblue}{\left(\color{black}{e^{4x}}\right)'}\] Recall the rule for differentiating an exponential of base e:\[\Large \frac{d}{dx}e^u\quad=\quad e^u\frac{du}{dx}\quad=\quad e^u\;(u)'\] Applying this rule gives us:\[\Large \color{black}{-5}\color{royalblue}{\left(\color{black}{e^{4x}}\right)'}\quad=\quad -5e^{4x}\color{royalblue}{\left(\color{black}{4x}\right)'}\]

zepdrix (zepdrix):

I took the derivative of the exponential ( it gave us the same thing back). Then we have to apply the chain rule, multiplying by the derivative of the inner function ( the contents of the exponent in this case ).

zepdrix (zepdrix):

Derivative of 4x?